OpenStudy (beccaboo333):

Write the equation of the line given by the following graph in slope-intercept form.

OpenStudy (beccaboo333):

im trying to attach the graph

OpenStudy (beccaboo333):

A. y = -2x + 86 B. y = -2x - 10 C. y = -2x - 8 D. y = -2x + 10